# On completeness of Foliated structures, and Null Killing fields

### v1

@@ 6.2. Incomplete examples with a null Killing field
in the basis $(T,X,V)$. The Euler-Arnold vector field in this case is
$w=\begin{pmatrix}t\\ x\\ v\end{pmatrix}\in{\mathfrak{g}}\mapsto ad^{*}_{w}w=\begin{pmatrix}-xv\\ 0\\ tv\end{pmatrix}$
@@
For $x=1$, we get that $\ddot{t}=-\dot{v}=-tv=t\dot{t}=(\frac{1}{2}t^{2})^{\prime}$. Hence, $\dot{t}=\frac{1}{2}t^{2}$. So $w(s)=(\frac{-2}{s},\frac{2}{s^{2}},1)^{T}$ is the velocity curve of an incomplete geodesic on the Lie group $G=\sf{Aff}(\mathbb{R})\times\mathbb{R}$ endowed with the left invariant metric coming from $q$.
@@
in the basis $(V,T,X,Y)$. In this case the Arnold-Euler vector field is
$w=\begin{pmatrix}t\\ x\\ y\\ v\end{pmatrix}\in{\mathfrak{g}}\mapsto ad^{*}_{w}w=\begin{pmatrix}-xv+y^{2}\\ 0\\ ty\\ tv\end{pmatrix}$
In other words we get the following system
@@
We see that the geodesic field is tangent to the hyperplane $y=0$, which is then totally geodesic. On the other hand, this hyperplane is the subalgebra ${\mathfrak{aff}}(\mathbb{R})\oplus\mathbb{R}$ with the (incomplete) Lorentzian scalar product given in Example 6.4 above. Hence incompleteness. We obtain a compact example by taking a quotient by a cocompact lattice.
@@ 7. Compact Lorentzian $3$-manifolds with an equicontinuous null Killing field
Let $(M,V)$ be a compact Lorentzian manifold of dimension $3$, with $V$ a lightlike Killing field. In this section, we study the geodesic completeness of $M$. As pointed out in the introduction, the situation is different, depending on the dynamics of $V$. The non-equicontinuous case being already known, we focus here on the equicontinuous case. In this case, the closure of the flow of $V$ in the isometry group is a torus. Denote it by $G$.
@@
An incomplete example is constructed in Paragraph 7.3.2, in the case where the flow of $V$ is periodic. The reader only needs Paragraph 7.2 to read it. Completeness results are given in special situations when the action of $V$ is not periodic, in paragraphs 7.3.3 and 7.3.1.
7.1. On the isometry group
We begin by a study of the connected component of the isometry group. We will see in particular that when it has dimension $\geq 3$, the manifold is flat.
Lemma 7.1.
Let $M$ be a connected compact Lorentzian $3$-manifold. If ${\sf{Isom}}(M)$ contains a $3$-dimensional connected abelian subgroup $H$, then $M$ is flat. Moreover, the identity component of the isometry group is isomorphic to $\mathbb{T}^{3}$.
Proof.
Assume first that $H$ is compact, i.e. $H=\mathbb{T}^{3}$. Let $K_{1},K_{2},K_{3}$ be three transversal Killing fields generated by the action of $H$. Let $C_{i}$ be the set of zeros of $K_{i}$, and let $C=\bigcup_{i=1}^{3}C_{i}$. Each of the $C_{i}$ has empty interior (this comes from the fact that the action of the isometry group is faithful). Consequently, $C$ is a closed subset of $M$, necessarily with empty interior (it is even of zero measure). So let $x\in M\smallsetminus C$. Then the $H$-orbit of $x$ is an open and closed subset of $M$, hence equal to $M$. This implies that the $H$-action is free and transitive, and $M$ is flat. Otherwise, $H$ is non-compact. By the construction above, there is a point $x_{0}\in M$ with a locally trivial isotropy. Hence, its orbit is a flat homogeneous open subset (of finite volume), we call it $U_{0}$. If $U_{0}$ is dense, then $M$ is flat, since the sectional curvature is zero on a dense subset of $M$. Otherwise, there is a point $x_{1}\in M\backslash U_{0}$, with also a locally trivial isotropy, and its orbit is a flat homogeneous orbit (of finite volume), we call it $U_{1}$. Moreover, $U_{0}\cap U_{1}=\emptyset$ and their volume is equal. This process must end for otherwise, $M$ would have infinite volume. When it ends, it means that there is some $U_{k}$ which is dense in the complement of $U_{0}\cup U_{1}\cup\cdots U_{k-1}$, so $M$ is covered by flat open subsets, hence flat. Now, since $M$ is flat (and compact), it is complete, and its isometry group is given by $\Gamma\backslash N_{\mathsf{Poi}^{1,2}}(\Gamma)$, where $N_{\mathsf{Poi}^{1,2}}(\Gamma)$ is the normalizer of $\Gamma:=\pi_{1}(M)$ in the Poincaré group. Since ${\sf{Isom}}(M)$ contains $H$, the normalizer of $\Gamma$ contains a $3$-dimensional (connected) abelian group, which is necessarily the translation group $\mathbb{R}^{3}$ by Observation 7.2 below. This group centralizes $\Gamma$, so $\Gamma$ is also made of pure translations (up to finite index). ∎
Observation 7.2.
The maximal dimension of an abelian connected subgroup of the Poincaré group ${\mathsf{Poi}}^{2,1}:={\sf{Isom}}({\sf{Mink}}^{1,2})$ is $3$. And it is identified with the pure translations.
@@ 7.1. On the isometry group
    If $\dim G=2$, then either $G$ coincides with the connected component of the isometry group, or, $M$ is flat and ${\sf{Isom}}^{\circ}(M)=\mathbb{T}^{3}$.
@@
$(2)$ Let ${\widehat{G}}:={\sf{Isom}}^{\circ}(M)$. First, recall that we have a bound on the dimension of ${\widehat{G}}$, namely $\dim{\widehat{G}}\leq 3+3$. The proof is based on the classification result on the connected component of the isometry group of a compact Lorentzian manifold [1, 2, 26].
Step 1: (Trivial non-compact part). Assume that the non-compact part of $\widehat{G}$ is non-trivial. We know via the classification result that the non-compact part (which, here, is either ${\sf{SL}}_{2}(\mathbb{R})\times\mathbb{R}^{k}$, ${\mathsf{Heis}}_{3}\times\mathbb{R}^{k}$, ${\mathsf{Osc}}\times\mathbb{R}^{k}$ or $\mathbb{R}^{k}$) commutes with the compact part, in particular with $\mathbb{T}^{2}$. Hence, ${\widehat{G}}$ contains a non-compact $3$-dimensional abelian subgroup given by the product of $\mathbb{T}^{2}$ with a one parameter group of the non-compact factor. This is absurd by Lemma 7.1.
Step 2: (Abelian compact part). Assume that the compact part of ${\widehat{G}}$ is non abelian. Since it contains $\mathbb{T}^{2}$, it is of the form $\mathbb{T}^{2}\times K$, where $K$ is, up to finite cover, isomorphic to either ${\sf{SO}}(3)$ or ${\sf{SO}}(3)\times\mathbb{S}^{1}$ (remember that $\dim K\leq 4$). Both cases will generate an abelian isotropy of dimension $\geq 2$, that represents (faithfully) in ${\sf{SO}}(1,2)$, which is absurd.
It follows from Step 1 and Step 2 that ${\widehat{G}}$ is compact abelian, isomorphic to $\mathbb{T}^{k}$, with $k\geq 2$. If $k=3$, $M$ is flat and ${\widehat{G}}=\mathbb{T}^{3}$ by Lemma 7.1. ∎
@@ 7.2. Existence of a $V$-invariant frame field
Let $(M,g,V)$ be an orientable compact Lorentzian manifold of dimension $3$, with $V$ a null Killing field. Then there is a null frame field $(V,X,Y)$ on $M$ which is invariant by the flow of $V$, and such that the metric in the basis $(V,Y,X)$ is given by $\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&1\\ \end{pmatrix}.$ Conversely, given a $V$-invariant frame field $(V,X,Y)$ in which the metric has the above form, $V$ is a null Killing field for this metric.
@@
The flow of $V$ preserves a Riemannian metric $h_{0}$ on $M$. Let $P_{0}$ (resp. $P$) be the $2$-dimensional distribution $h_{0}$-orthogonal ($g$-orthogonal) to $V$. Then $P_{0}\cap P$ is a line bundle over $M$. Define a vector field $X$ on $M$ tangent to the line bundle, and satisfying $g(X,X)=1$. Let $Y$ be the isotropic vector field on $M$ which is $g$-orthogonal to $X$ and satisfies $g(V,Y)=1$. Since the flow of $V$ is isometric for both $g$ and $h_{0}$, these vector fields are invariant by the flow of $V$, i.e. $[V,X]=[V,Y]=0$. For the converse, one proves that $\nabla V$ is skew-symmetric for $g$, using that the scalar products $g(U,W)$ are $V$-invariant for $U,W\in\{V,X,Y\}$. ∎
In particular, the foliation tangent to $V$ is trasnversally parallelizable. The action of $V$ is assumed to be equicontinuous. So the closure of the leaves of $V$ are the fibers of a fibration defined by the (proper) action of $G$ (the closure of the flow of $V$ in the isometry group). When the action of $V$ is free (observe that the action also preserves the orientation of $M$), the base space of this fibration is a compact and orientable manifold. Its Euler characteristic is zero, since the $V$-action preserves non-vanishing vector fields transversal to $V$, who then project to well defined vector fields on the base space. The base is therefore diffeomorphic to a torus (of dimension $\leq 2$), which we denote by $\Sigma$.
@@
In particular, $V$ is parallel if and only if $f=0$. Moreover, $f$, $h$ and $\mu$ are constant along the orbits of $V$, so they project to $\Sigma$.

### v2

@@ 6.2. Incomplete examples with a null Killing field
in the basis $(T,X,V)$. Since $V$ is central, $\mathrm{ad}_{V}$ is skew-symmetric with respect to $q$. Therefore, the left invariant vector field generated by $V$ is null and Killing. The Euler-Arnold vector field in this case is
$w=\begin{pmatrix}t\\ x\\ v\end{pmatrix}\in{\mathfrak{g}}\mapsto\mathrm{ad}^{*}_{w}w=\begin{pmatrix}-xv\\ 0\\ tv\end{pmatrix}.$
@@
Moreover, $q(\omega,\omega)=t^{2}+2xv=C$, $C\in\mathbb{R}$. For $x=1$, we get that $\ddot{t}=-\dot{v}=-tv=t\dot{t}=(\frac{1}{2}t^{2})^{\prime}$. Hence $\dot{t}=\frac{1}{2}(t^{2}-C)$. So $w(s)=(\frac{-2}{s},\frac{2}{s^{2}},1)^{T}$ is the velocity curve of an incomplete lightlike geodesic on the Lie group $G=\sf{Aff}(\mathbb{R})\times\mathbb{R}$ endowed with the left invariant metric coming from $q$.
@@
in the basis $(T,X,Y,V)$. As in the previous example, the left invariant vector field generated by $V$ is null and Killing. In this case, the Arnold-Euler vector field is
$w=\begin{pmatrix}t\\ x\\ y\\ v\end{pmatrix}\in{\mathfrak{g}}\mapsto\mathrm{ad}^{*}_{w}w=\begin{pmatrix}-xv+y^{2}\\ 0\\ ty\\ tv\end{pmatrix}.$
In other words, we get the following system
@@
We see that the geodesic field is tangent to the hyperplane $y=0$, which is then totally geodesic. On the other hand, this hyperplane is the subalgebra ${\mathfrak{aff}}(\mathbb{R})\oplus\mathbb{R}$ with the (incomplete) Lorentzian scalar product given in Example 6.4 above. Hence incompleteness. We obtain a locally homogeneous compact example by taking the quotient of ${\mathsf{Sol}}\times\mathbb{R}$ by a cocompact lattice.
@@ 7. Compact Lorentzian $3$-manifolds with an equicontinuous null Killing field
Let $(M,V)$ be a compact Lorentzian manifold of dimension $3$, with $V$ a null Killing field. In this section, we study the geodesic completeness of $M$. As pointed out in the introduction, the situation is different, depending on the dynamics of $V$. The non-equicontinuous case being already known in dimension $3$, we focus on the equicontinuous case. In this case, the closure of the flow of $V$ in the isometry group is a torus. Denote it by $G$.
@@
We begin by a study of the connected component of the isometry group in Paragraph 7.1. We will see in particular that when it has dimension $\geq 3$, the manifold is flat, hence complete. An incomplete example is constructed in Paragraph 7.3.2, in the case where the flow of $V$ is periodic. The reader can proceed directly to Paragraph 7.2 for this example (which is independent of Paragraph 7.1). Completeness results are also given in Paragraphs 7.3.3 and 7.3.1, in special situations where the action of $V$ is not periodic.
7.1. Isometry groups
The aim of this section is to prove the following result
@@ 7.1. Isometry groups
    If $\dim G=2$, then either the connected component of the isometry group is $G\simeq\mathbb{T}^{2}$, or, $M$ is flat and ${\sf{Isom}}^{o}(M)=\mathbb{T}^{3}$.
We start with the following two lemmas
Lemma 7.2.
Let $M$ be a connected compact Lorentzian $3$-manifold. If ${\sf{Isom}}(M)$ contains a $3$-dimensional connected solvable subgroup $H$, then $M$ is $H$-homogeneous. If $H$ is abelian, then $M$ is flat, and the identity component of the isometry group is isomorphic to $\mathbb{T}^{3}$.
@@
Assume first that $H$ is compact, i.e. $H=\mathbb{T}^{3}$. Let $K_{1},K_{2},K_{3}$ be three transversal Killing fields generated by the action of $H$. Let $C_{i}$ be the set of zeros of $K_{i}$, and let $C=\bigcup_{i=1}^{3}C_{i}$. Each of the $C_{i}$ has empty interior (this comes from the fact that the action of the isometry group is faithful). Consequently, $C$ is a closed subset of $M$ with empty interior. So let $x\in M\smallsetminus C$. Then the $H$-orbit of $x$ is an open and closed subset of $M$, hence equal to $M$. This implies that the $H$-action is locally free and transitive. Now, assume that $H$ is non-compact. By the construction above, there exists a point $x_{0}\in M$ with trivial local isotropy. Hence, its orbit is a homogeneous open subset of $M$ (of finite volume). Since the action of $H$ on $U_{0}$ is locally free, this orbit is given by $U_{0}=H/\Gamma_{0}$, where the isotropy group $\Gamma_{0}$ is a discrete subgroup of $H$. Now, since $H/\Gamma_{0}$ has finite volume (with respect to the volume element induced by the Lorentzian metric of $U_{0}$), and $H$ is solvable, the subgroup $\Gamma_{0}$ must be a cocompact lattice in $H$. Therefore, $U_{0}$ is compact, making it a closed subset of $M$. Finally, since $M$ is connected, we conclude that $U_{0}=M$. This proves the first part of the lemma. Now, if $H$ is abelian, then $M$ is clearly flat. In this case, $M$ is flat and compact, hence complete. So its full isometry group is given by $\Gamma\backslash N(\Gamma)$, where $N(\Gamma)$ is the normalizer of $\Gamma:=\pi_{1}(M)$ in the Poincaré group $\mathsf{Poi}^{2,1}={\sf{O}}(2,1)\ltimes\mathbb{R}^{2+1}$. As $H\subset{\sf{Isom}}(M)$, the normalizer $N(\Gamma)$ contains a $3$-dimensional connected abelian group, which must be the translation group $\mathbb{R}^{3}$ by Observation 7.3 below. Hence, the subgroup of translations $\mathbb{R}^{3}$ centralizes $\Gamma$, and this implies that $\Gamma$ consists of pure translations. ∎
Observation 7.3.
The maximal dimension of an abelian connected subgroup of the Poincaré group ${\mathsf{Poi}}^{2,1}:={\sf{Isom}}({\sf{Mink}}^{2,1})$ is $3$. Moreover, it is identified with the subgroup of pure translations.
Lemma 7.4.
A compact Lorentzian $3$-manifold of constant curvature has an isometry group of dimension $\leq 3$.
Proof.
In this case, $M$ is complete by [6, 21], hence can be identified with a compact quotient of ${\sf{Mink}}^{2,1}$ or $\widetilde{{\sf{AdS}}^{3}}$. In both cases $M$ is locally homogeneous. Thus, the stabilizer in ${\sf{Isom}}(M)$ of a point $p\in M$ is isomorphic to the stabilizer of any other point.
We start with the flat case. In this case, $M=\Gamma\backslash{\sf{Mink}}^{2,1}$, where $\Gamma\subset{\sf{O}}(2,1)\ltimes\mathbb{R}^{3}$ is a discrete subgroup that acts properly, freely, and cocompactly. Assume $\dim{\sf{Isom}}(M)\geq 4$. We will show that $M$ is homogeneous under the action of the isometry group. Let $I$ denote the isotropy of some point in $M$. If $\dim I=1$, then the orbits of the isometry group are all open, so all closed. Since $M$ is connected, it consists of a single orbit, so homogeneity follows. Now, if $\dim I\geq 2$, then, by [3, Proposition 2.3], it is, in fact, the whole ${\sf{O}}(2,1)$ up to finite index. On the other hand, the isometry group of $M$ is given by $\Gamma\backslash N(\Gamma)$, where $N(\Gamma)$ is the normalizer of $\Gamma$ in ${\sf{O}}(2,1)\ltimes\mathbb{R}^{3}$. In particular, the isotropy group ${\sf{O}}(2,1)$ normalizes $\Gamma$. This is impossible, since such a group $\Gamma$ must be finite. Finally, $M$ is homogeneous as claimed. In the latter case, $M$ is necessarily the quotient of $\mathbb{R}^{3}$ by a lattice $\Gamma$ of translations. This can be shown by elementary arguments (in dimension $3$), but we refer to [4, Corollary 4.9] for this fact. The centralizer of $\Gamma$ is $\mathbb{R}^{3}$, so the identity component of the isometry group is isomorphic to $\mathbb{T}^{3}$.
In the anti-de Sitter case, $M$ is a compact quotient $\widetilde{\Gamma}\backslash\widetilde{{\sf{AdS}}^{3}}$ by a discrete subgroup $\widetilde{\Gamma}\subset{\sf{Isom}}(\widetilde{{\sf{AdS}}^{3}})$ acting properly, freely and cocompactly. Assume $\dim{\sf{Isom}}(M)\geq 4$. Then, the normalizer of $\widetilde{\Gamma}$ in ${\sf{Isom}}(\widetilde{{\sf{AdS}}^{3}})$ has dimension at least $4$. Denote by $Z$ the center of ${\sf{Isom}}(\widetilde{{\sf{AdS}}^{3}})$, it is isomorphic to $\mathbb{Z}$, and ${\sf{Isom}}(\widetilde{{\sf{AdS}}^{3}})/Z={\sf{Isom}}({\sf{AdS}}^{3})$. Denote by $\pi:{\sf{Isom}}(\widetilde{{\sf{AdS}}^{3}})\to{\sf{Isom}}({\sf{AdS}}^{3})$ the natural projection. A covering of $M$ also has an isometry group of dimension $\geq 4$. So, up to taking a finite cover, we can assume that $\widetilde{\Gamma}\subset{\sf{Isom}}^{\mathsf{o}}(\widetilde{{\sf{AdS}}^{3}})$. Then, by [22, Theorem 7.2], $\widetilde{\Gamma}$ intersects the center $Z$ non-trivially, which guarantees that the quotient group $\Gamma:=\widetilde{\Gamma}/\widetilde{\Gamma}\cap Z$ is a subgroup of ${\sf{Isom}}({\sf{AdS}}^{3})$ that acts properly, freely and cocompactly on ${\sf{AdS}}^{3}:=\widetilde{{\sf{AdS}}^{3}}/Z$. The manifold $M$ is a (finite) cover of $\overline{M}:=\Gamma\backslash{\sf{AdS}}^{3}=\widetilde{\Gamma}Z\backslash\widetilde{{\sf{AdS}}^{3}}$. By assumption, $\widetilde{\Gamma}$ has a normalizer, denoted by $C$, of dimension at least $4$ in ${\sf{Isom}}(\widetilde{{\sf{AdS}}^{3}})$. Now, the fact that the kernel of the projection $\pi$ is discrete implies that $\Gamma$ has a normalizer of dimension at least $4$ in ${\sf{Isom}}({\sf{AdS}}^{3})$, given by the projection of $C$. The advantage of considering $\overline{M}$ is that we have a better understanding of its algebraic form. Indeed, by [22, Theorem 5.2], up to conjugacy, there exists a (torsion free) cocompact lattice $\Gamma_{0}$ in ${\sf{PSL}}_{2}(\mathbb{R})$, and a representation $\rho:\Gamma_{0}\to{\sf{PSL}}_{2}(\mathbb{R})$, such that $\Gamma=\{(x,\rho(x)),x\in\Gamma_{0}\}$. Let $p_{1}:{\sf{PSL}}_{2}(\mathbb{R})\times{\sf{PSL}}_{2}(\mathbb{R})\to{\sf{PSL}}_{2}(\mathbb{R})$ be the projection on the first factor. Since $\dim C\geq 4$, $\Gamma_{0}$ has at least a $1$-dimensional normalizer in ${\sf{PSL}}_{2}(\mathbb{R})$, given by $p_{1}(C)$. This normalizer acts on the compact hyperbolic surface $\Gamma_{0}\backslash\mathbb{H}^{2}$ by (at least) a $1$-dimensional group of isometries, contradicting the fact that the isometry group of a compact hyperbolic surface is finite. ∎
Recall that the foliation tangent to $V^{\perp}$ is lightlike and totally geodesic. For the proof of Proposition 7.1, we need the following general fact
Fact 7.5 (Theorem 11, [38]).
Let $M$ be a compact Lorentzian $3$-manifold with a $C^{0}$ lightlike geodesic foliation. Then the universal cover of $M$ is diffeomorphic to $\mathbb{R}^{3}$.
Proof of Proposition 7.1.
@@
$(2)$ Let ${\widehat{G}}:={\sf{Isom}}^{\mathsf{o}}(M)$. First, recall that we have a bound on the dimension of ${\widehat{G}}$, namely, $\dim{\widehat{G}}\leq 6$. The proof is based on the classification result on the connected component of the isometry group of a compact Lorentzian manifold [1, 2, 37]. By [37, Theorem 1.5], the Lie algebra of ${\widehat{G}}$ decomposes as $\widehat{g}=\mathbb{R}^{k}\oplus\mathcal{K}\oplus\mathcal{H}$, where $\mathcal{K}$ is the Lie algebra of a compact semisimple Lie group and $\mathcal{H}$ is isomorphic to either ${\mathfrak{sl}}_{2}(\mathbb{R}),{\mathfrak{heis}}$, or ${\mathfrak{osc}}$ the Lie algebra of an oscillator group. In our case, $\mathbb{R}^{2}\subset\widehat{g}$. We need to consider the following two cases:
  • I.
    If $\dim(\widehat{g})=3$, then $\widehat{g}$ is isomorphic to either $\mathbb{R}^{3}$ or ${\mathfrak{heis}}_{3}$. Indeed, the ${\mathfrak{sl}}_{2}(\mathbb{R})$ and ${\mathfrak{so}}(3)$ cases are excluded, since their maximal abelian subgroup has dimension $1$. Thus, $\widehat{G}$ is solvable, and by Lemma 7.2, $M=\widehat{G}/\Gamma$, where $\Gamma$ is a cocompact lattice in $\widehat{G}$. Since ${\widehat{G}}$ contains $\mathbb{T}^{2}$, the normalizer of $\Gamma$ in ${\widehat{G}}$ has dimension $\geq 2$. Therefore, $\widehat{g}=\mathbb{R}^{3}$. In this case, $M$ is flat by Lemma 7.2.
  • II.
    If $\dim(\widehat{g})\geq 4$ then $M$ is locally homogeneous by [8, Theorem 4.1]. Let $I$ be the isotropy group of some point. The case where $\dim I=2$ is not possible by [3, Proposition 2.3]. We are left with two situations. If $\dim I\geq 3$, then $M$ has constant curvature. So $M$ is a compact constant curvature Lorentzian $3$-manifold with an isometry group of dimension $\geq 4$. This is not possible by Lemma 7.4. If $\dim I=1$, the orbits of $\widehat{G}$ in $M$ are all open, and therefore also closed. Since $M$ is connected, it consists of a single orbit, meaning $M$ is $\widehat{G}$-homogeneous. Consequently, $M$ can be identified with ${\widehat{G}}/I$, where $I$ is the isotropy of some point. In particular, $\dim\widehat{g}=4$. In this case, the possibilities for $\widehat{g}$ are $\mathbb{R}^{4}$, $\mathbb{R}\oplus{\mathfrak{sl}}_{2}(\mathbb{R})$, $\mathbb{R}\oplus{\mathfrak{heis}}_{3}$, $\mathbb{R}\oplus{\mathfrak{so}}(3)$, and ${\mathfrak{osc}}$. By Lemma 7.2, we can rule out all possibilities except for $\mathbb{R}\oplus{\mathfrak{so}}(3)$. In the latter case, ${\widehat{G}}$ is isomorphic to ×$\mathbb{R}\times K$ or $\mathbb{S}^{1}\times K$, where $K$ is finitely covered by $\mathbb{S}^{3}$. In both cases, $M$ has a non-contractible universal cover, diffeomorphic to either $\mathbb{S}^{3}$ or ×$\mathbb{R}\times\mathbb{S}^{2}$. This is not possible by Fact 7.5.
Finally, either $\dim{\widehat{G}}=3$, and $M$ is flat. Or $\dim{\widehat{G}}=2$, in which case ${\widehat{G}}=G$. This ends the proof. ∎
@@ 7.2. Existence of a $V$-invariant frame field
Let $(M,g,V)$ be an orientable compact Lorentzian manifold of dimension $3$, with $V$ an equicontinuous null Killing field. Then, there exists a null frame field $(V,X,Y)$ on $M$ which is invariant by the flow of $V$, and such that the metric in the basis $(V,Y,X)$ is given by $\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&1\\ \end{pmatrix}.$ Conversely, given a $V$-invariant frame field $(V,X,Y)$ in which the metric has the above form, $V$ is a null Killing field for this metric.
@@
The flow of $V$ preserves a Riemannian metric $h_{0}$ on $M$. Let $P_{0}$ (resp. $P$) be the $2$-dimensional distribution $h_{0}$-orthogonal (resp. $g$-orthogonal) to $V$. Then $P_{0}\cap P$ is a line bundle over $M$. Up to double cover, we can define a vector field $X$ on $M$ tangent to the line bundle, and satisfying $g(X,X)=1$. Let $Y$ be the null vector field on $M$ which is $g$-orthogonal to $X$ and satisfies $g(V,Y)=1$. Since the flow of $V$ is isometric for both $g$ and $h_{0}$, these vector fields are invariant by the flow of $V$, i.e. $[V,X]=[V,Y]=0$. For the converse, one proves that $\nabla V$ is skew-symmetric for $g$, using that the scalar products $g(U,W)$ are $V$-invariant for $U,W\in\{V,X,Y\}$. ∎
In particular, the foliation tangent to $V$ is trasnversally parallelizable. The action of $V$ is assumed to be equicontinuous. So the closure of the leaves of $V$ are the fibers of a fibration defined by the (proper) action of $G$, the closure of the flow of $V$ in the isometry group. When the action of $V$ is free (note that the action also preserves the orientation of $M$), the base space of this fibration is a compact, orientable manifold. Its Euler characteristic is zero, as the $V$-action preserves non-vanishing vector fields transverse to $V$, which project to well-defined vector fields on the base space. Therefore, the base is diffeomorphic to a torus (of dimension $\leq 2$), which we denote by $\Sigma$.
@@
In particular, $V$ is parallel if and only if $f=0$. Moreover, since $f$, $h$ and $\mu$ are constant along the orbits of $V$, they project to $\Sigma$.